# Kernelized Concept Erasure

### v5

@@ Abstract
The representation space of neural models for textual data emerges in an unsupervised manner during training. Understanding how those representations encode human-interpretable concepts is a fundamental problem. One prominent approach for the identification of concepts in neural representations is searching for a linear subspace whose erasure prevents the prediction of the concept from the representations. However, while many linear erasure algorithms are tractable and interpretable, neural networks do not necessarily represent concepts in a linear manner. To identify non-linearly encoded concepts, we propose a kernelization of a linear minimax game for concept erasure. We demonstrate that it is possible to prevent specific non-linear adversaries from predicting the concept. However, the protection does not transfer to different non-linear adversaries. Therefore, exhaustively erasing a non-linearly encoded concept remains an open problem.
@@ 1 Introduction
Concept erasure is tightly related to concept identification. Once we have successfully removed a given concept, e.g., color, from a representation $\mathbf{x}_{n}$, it is reasonable to argue that the erasure function $r$ has meaningfully identified the concept within the representation space. In the rest of the paper, we will say that $r(\cdot)$ neutralizes the concept in the representation space. For instance, we say that the $r$ in our example neutralizes the concept of a laptop’s color. It follows that concept identification is related to bias mitigation Bolukbasi et al. (2016); Gonen and Goldberg (2019); Maudslay et al. (2019), e.g., one may want to identify and remove the gender encoded in learned word representations produced by an embedding method such as word2vec (Mikolov et al., 2013) or GloVe (Pennington et al., 2014). Indeed, the empirical portion of this paper will focus on removing gender from word representations in order to mitigate bias.
Previous work on concept erasure (Ravfogel et al., 2021) focuses on the linear case, i.e., where $r$ is a linear function. While linear concept erasure methods have certainly found success Bolukbasi et al. (2016), there is no a-priori reason to suspect that neural networks encode concepts in a linear manner. In this work, we take the first step toward the goal of identifying a non-linear function $r(\cdot)$ and a corresponding non-linearly encoded concept subspace. We directly build on Ravfogel et al. (2022), who cast linear concept erasure as a minimax game. Under their formulation, the function $r(\cdot)$ learns to remove the concept, while an adversary tries to predict the concept. We extend their work by deriving a class of general minimax games based on kernelization that largely maintains the tractability of the linear approach. Our kernelized method performs concept erasure in a reproducing kernel Hilbert space, which may have a much higher dimensionality (Schölkopf and Smola, 2002) and correspond to a non-linear subspace of the original representation space.
Empirically, we experiment with gender erasure from GloVe and BERT representations. We show that a kernelized adversary can classify the gender of the representations with over 99% accuracy if $r(\cdot)$ is taken to be a linear function. This gives us concrete evidence that gender is indeed encoded non-linearly in the representations. We further find that solving our kernelized minimax game yields an erasure function $r(\cdot)$ that protects against an adversary that shares the same kernel. However, we also find that it is difficult to protect against all kernelized adversaries at once: Information removed by one kernel type can be recovered by adversaries using other kernel types. That is, the gender concept is not exclusively encoded in a space that corresponds to any one kernel. This suggests that non-linear concept erasure is very much an open problem.
@@ 2 Linear Concept Erasure
We provide an overview of the linear minimax formulation before we introduce its kernelization.
@@ 2.1 Notation
Let ${\mathcal{D}}=\{(y_{n},\mathbf{x}_{n})\}_{n=1}^{N}$ be a dataset of $N$ concept–representation pairs, where the labels $y_{n}$ represent the concept to be neutralized. The goal of linear concept erasure is to learn a linear erasure function $r(\cdot)$ from ${\mathcal{D}}$ such that it is impossible to predict $y_{n}$ from the modified representations $r(\mathbf{x}_{n})$. We focus on classification, where the concept labels $y_{n}$ are derived from a finite set $\{1,\dots,V\}$ of $V$ discrete values, and the representations $\mathbf{x}_{n}\in\mathbb{R}^{D}$ are $D$-dimensional real column vectors. To predict the concepts labels $y_{n}$ from the representations $\mathbf{x}_{n}$, we make use of (and later kernelize) classifiers that are linear models, i.e., classifiers of the form $\widehat{y}_{n}={\boldsymbol{\theta}}^{\top}\mathbf{x}_{n}$ where ${\boldsymbol{\theta}}\in\Theta\subseteq\mathbb{R}^{D}$ is a column vector of parameters that lives in a compact set $\Theta$. We also consider an arbitrary loss function $\ell(\cdot,\cdot)\geq 0$ where $\ell(y_{n},\widehat{y}_{n})$ tells us how close the prediction $\widehat{y}_{n}$ is to $y_{n}$.
@@
Using this notation, linear concept erasure is realized by identifying a linear concept subspace whose neutralization, achieved through an orthogonal projection matrix, prevents the classifier from predicting the concept. We define $\mathcal{P}_{k}$ as the set of all $D\times D$ orthogonal projection matrices that neutralize a rank $k$ subspace. More formally, we have that $P\in\mathcal{P}_{k}\leftrightarrow P=I_{D}-W^{\top}W,W\in\mathbb{R}^{k\times D},WW^{\top}=I_{k}$, where $I_{k}$ denotes the $k\times k$ identity matrix and $I_{D}$ denotes the $D\times D$ identity matrix. We say that the matrix $P$ neutralizes the $k$-dimensional rowspace of $W$.
2.2 A Linear Minimax Game
@@
Following this formalization, it is natural to define a minimax game (Neumann and Morgenstern, 1944) between a projection matrix $P\in\mathcal{P}_{k}$ that aims to remove the concept subspace, and a linear model parameterized by ${\boldsymbol{\theta}}$ that aims to recover it:
$\min_{{\boldsymbol{\theta}}\in\Theta}\,\max_{{P\in\mathcal{P}_{k}}}\,\sum_{n=1}^{N}\ell\Big{(}y_{n},\,{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}\Big{)}$ (1)
This is a special case of the general adversarial framework (Goodfellow et al., 2014). However, in our case, the predictor ${\boldsymbol{\theta}}$ does not interact with the original input $\mathbf{x}_{n}$. Instead, the classifier attempts to predict the concept label $y_{n}$ from $P\mathbf{x}_{n}$. We now give a concrete example of the linear concept: When we have a binary logistic loss ($V=2$) and a $k$-dimensional neutralized subspace, the game takes the following form:
$\min_{{\boldsymbol{\theta}}\in\mathbb{R}^{D}}$ $\max_{P\in\mathcal{P}_{k}}\,\,-\Bigg{(}\sum_{n=1}^{N}y_{n}\log\frac{\exp{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}}{1+\exp{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}}$ (2)
$+\Big{(}1-y_{n}\Big{)}\log\frac{1}{1+\exp{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}}\Bigg{)}$
In this paper, we focus on $k=1$, i.e., we aim to identify a 1-dimensional concept subspace. [^2] This construction can be extended to subspaces of dimension $k>1$ using a version of the Gram–Schmidt process in the RKHS. The optimization over $\mathcal{P}_{k}$ in Eq. 1 renders the game non-convex. To address this issue, Ravfogel et al. (2022) propose a concave relaxation of the objective:
@@
$\min_{{\boldsymbol{\theta}}\in\Theta}\,\max_{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}P\in\mathcal{F}_{k}}}\,\sum_{n=1}^{N}\ell\Big{(}y_{n},\,{\boldsymbol{\theta}}^{\top}P\mathbf{x}_{n}\Big{)}$ (3)
where the relaxation is shown in gray. In words, instead of optimizing over rank-$k$ projection matrices, a non-convex set, we optimize over its convex hull, the Fantope (Boyd and Vandenberghe, 2014):
$\mathcal{F}_{k}=\{A\in\mathcal{S}^{D}\mid I_{D}\succcurlyeq A\succcurlyeq 0,\mathrm{tr}(A)=k\}$ (4)
@@ 3 Non-linear Concept Erasure
The underlying assumption motivating kernel methods is that the features needed for the task live in a reproducing kernel Hilbert space (RKHS; Aronszajn, 1950). At an intuitive level, an RKHS allows us to extend some of results from linear algebra to potentially infinite-dimensional vector spaces (Canu and Smola, 2006). The main technical contribution of this paper is the derivation of a kernelized version of the linear adversarial game presented in Eq. 3. We perform this derivation in this section after providing some background on kernel methods. We show that the resulting kernelized game is both non-convex and too computationally heavy to solve directly. Thus, we introduce a Nyström approximation that results in an efficient and light-weight formulation, given in Eq. 10, that is still able to isolate concepts encoded non-linearly.

### v6

@@ Abstract
The representation space of neural models for textual data emerges in an unsupervised manner during training. Understanding how those representations encode human-interpretable concepts is a fundamental problem. One prominent approach for the identification of concepts in neural representations is searching for a linear subspace whose erasure prevents the prediction of the concept from the representations. However, while many linear erasure algorithms are tractable and interpretable, neural networks do not necessarily represent concepts in a linear manner. To identify non-linearly encoded concepts, we propose a kernelization of a linear maximin game for concept erasure. We demonstrate that it is possible to prevent specific non-linear adversaries from predicting the concept. However, the protection does not transfer to different non-linear adversaries. Therefore, exhaustively erasing a non-linearly encoded concept remains an open problem.
@@ 1 Introduction
Concept erasure is tightly related to concept identification. Once we have successfully removed a given concept, e.g., color, from a representation $\boldsymbol{x}_{n}$, it is reasonable to argue that the erasure function $r$ has meaningfully identified the concept within the representation space. In the rest of the paper, we will say that $r(\cdot)$ neutralizes the concept in the representation space. For instance, we say that the $r$ in our example neutralizes the concept of a laptop’s color. It follows that concept identification is related to bias mitigation Bolukbasi et al. (2016); Gonen and Goldberg (2019); Maudslay et al. (2019), e.g., one may want to identify and remove the gender encoded in learned word representations produced by an embedding method such as word2vec (Mikolov et al., 2013) or GloVe (Pennington et al., 2014). Indeed, the empirical portion of this paper will focus on removing gender from word representations in order to mitigate bias.
Previous work on concept erasure (Ravfogel et al., 2021) focuses on the linear case, i.e., where $r$ is a linear function. While linear concept erasure methods have certainly found success Bolukbasi et al. (2016), there is no a-priori reason to suspect that neural networks encode concepts in a linear manner. In this work, we take the first step toward the goal of identifying a non-linear function $r(\cdot)$ and a corresponding non-linearly encoded concept subspace. We directly build on Ravfogel et al. (2022), who cast linear concept erasure as a maximin game. Under their formulation, the function $r(\cdot)$ learns to remove the concept, while an adversary tries to predict the concept. We extend their work by deriving a class of general maximin games based on kernelization that largely maintains the tractability of the linear approach. Our kernelized method performs concept erasure in a reproducing kernel Hilbert space, which may have a much higher dimensionality (Schölkopf and Smola, 2002) and correspond to a non-linear subspace of the original representation space.
Empirically, we experiment with gender erasure from GloVe and BERT representations. We show that a kernelized adversary can classify the gender of the representations with over 99% accuracy if $r(\cdot)$ is taken to be a linear function. This gives us concrete evidence that gender is indeed encoded non-linearly in the representations. We further find that solving our kernelized maximin game yields an erasure function $r(\cdot)$ that protects against an adversary that shares the same kernel. However, we also find that it is difficult to protect against all kernelized adversaries at once: Information removed by one kernel type can be recovered by adversaries using other kernel types. That is, the gender concept is not exclusively encoded in a space that corresponds to any one kernel. This suggests that non-linear concept erasure is very much an open problem.
@@ 2 Linear Concept Erasure
We provide an overview of the linear maximin formulation before we introduce its kernelization.
@@ 2.1 Notation
Let ${\mathcal{D}}=\{(\boldsymbol{x}_{n},y_{n})\}_{n=1}^{N}$ be a dataset of $N$ concept–representation pairs, where the labels $y_{n}$ represent the concept to be neutralized. The goal of linear concept erasure is to learn a linear erasure function $r(\cdot)$ from ${\mathcal{D}}$ such that it is impossible to predict $y_{n}$ from the modified representations $r(\boldsymbol{x}_{n})$. We focus on classification, where the concept labels $y_{n}$ are derived from a finite set $\{1,\dots,V\}$ of $V$ discrete values, and the representations $\boldsymbol{x}_{n}\in\mathbb{R}^{D}$ are $D$-dimensional real column vectors. To predict the concepts labels $y_{n}$ from the representations $\boldsymbol{x}_{n}$, we make use of (and later kernelize) classifiers that are linear models, i.e., classifiers of the form $\widehat{y}_{n}={\boldsymbol{\theta}}^{\top}\boldsymbol{x}_{n}$ where ${\boldsymbol{\theta}}\in\Theta\subseteq\mathbb{R}^{D}$ is a column vector of parameters that lives in a compact set $\Theta$. We also consider an arbitrary loss function ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}(\cdot,\cdot)\geq 0$ where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}(y_{n},\widehat{y}_{n})$ tells us how close the prediction $\widehat{y}_{n}$ is to $y_{n}$.
@@
Using this notation, linear concept erasure is realized by identifying a linear concept subspace whose neutralization, achieved through an orthogonal projection matrix, prevents the classifier from predicting the concept. We define $\mathcal{P}_{k}$ as the set of all $D\times D$ orthogonal projection matrices that neutralize a rank $k$ subspace. More formally, we have that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\in\mathcal{P}_{k}\iff{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{I}}_{D}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{W}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{W}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{W}}\in\mathbb{R}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}\times D},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{W}}^{\top}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{I}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}$, where ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{I}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}}$ denotes the ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}$ identity matrix and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{I}}_{D}$ denotes the $D\times D$ identity matrix. We say that the matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ neutralizes the ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}$-dimensional rowspace of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{W}}$.
2.2 A Linear Maximin Game
@@
Following this formalization, it is natural to define a maximin game (Neumann and Morgenstern, 1944) between a projection matrix that aims to remove the concept subspace, and a linear model parameterized by ${\boldsymbol{\theta}}$ that aims to recover it:
@@
$\max_{{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\in\mathcal{P}_{k}}}\,\min_{{\boldsymbol{\theta}}\in\Theta}\,\sum_{n=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}\Big{(}y_{n},\,{\boldsymbol{\theta}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{x}_{n}\Big{)}.$ (1)
Eq. 1 is a sequential game. The first player chooses an orthogonal projection matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ to hinder the prediction of $y_{n}$ from $\boldsymbol{x}_{n}$. In response, the second player chooses the parameters of a linear model ${\boldsymbol{\theta}}$, with knowledge of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$, to best predict $y_{n}$ from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{x}_{n}$. This is a special case of the general adversarial framework (Goodfellow et al., 2014). However, in our case, the predictor ${\boldsymbol{\theta}}$ does not interact with the original input $\boldsymbol{x}_{n}$. Instead, the classifier attempts to predict the concept label $y_{n}$ from ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{x}_{n}$. We now give a concrete example of the linear concept: When we have a binary logistic loss ($V=2$) and a ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}$-dimensional neutralized subspace, the game takes the following form:
$\max_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\in\mathcal{P}_{k}}$ $\,\min_{{\boldsymbol{\theta}}\in\Theta}\,\,-\Bigg{(}\sum_{n=1}^{N}y_{n}\log\frac{\exp{\boldsymbol{\theta}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{x}_{n}}{1+\exp{\boldsymbol{\theta}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{x}_{n}}$ (2)
$+\Big{(}1-y_{n}\Big{)}\log\frac{1}{1+\exp{\boldsymbol{\theta}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{x}_{n}}\Bigg{)}.$
In this paper, we focus on ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}=1$, i.e., we aim to identify a 1-dimensional concept subspace. [^2] This construction can be extended to subspaces of dimension $k>1$ using a version of the Gram–Schmidt process in the RKHS. Every $D\times D$ orthogonal projection matrix of rank $D-1$ can be written as follows
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}^{\bot}_{\boldsymbol{w}}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{I}}-\frac{\boldsymbol{w}\boldsymbol{w}^{\top}}{\boldsymbol{w}^{\top}\boldsymbol{w}},$ (3)
itself a differentiable function of $\boldsymbol{w}$. This results in the following game:
$\!\!\!\!\max_{\boldsymbol{w}\in\mathbb{R}^{D}}\min_{{\boldsymbol{\theta}}\in\Theta}\,\sum_{n=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}\left(y_{n},\,{\boldsymbol{\theta}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}^{\bot}_{\boldsymbol{w}}\boldsymbol{x}_{n}\right).$ (4)
author=ryan,color=violet!40,size=,fancyline,caption=,]Why did we switch form ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}^{\top}$
@@ 3 Non-linear Concept Erasure
The underlying assumption motivating kernel methods is that the features needed for the task live in a reproducing kernel Hilbert space (RKHS; Aronszajn, 1950). At an intuitive level, an RKHS allows us to extend some of results from linear algebra to potentially infinite-dimensional inner prodcut spaces (Canu and Smola, 2006). The main technical contribution of this paper is the derivation of a kernelized version of the linear adversarial game presented in Eq. 4. We perform this derivation in this section after providing some background on kernel methods. We show that the resulting kernelized game is both non-convex and too computationally heavy to solve directly. Thus, we introduce a Nyström approximation that results in an efficient and light-weight formulation, given in Eq. 10, that is still able to isolate concepts encoded non-linearly.